Circular No. 6017

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) R. M. West, European Southern Observatory, has measured the nuclei

on CCD images obtained during July 1–8 by O. Hainaut, R. Schulz, M. Carollo, C. Alard and A. Cimatti with the 3.5-m New Technology Telescope and 1.5-m Danish telescope. Reductions were with the help of southern Sky Atlas plates and provisional Hipparcos reference-star positions provided by M. Perryman and C. Turon. S. Nakano, Sumoto, Japan, has computed improved orbits and the following times (corrected for light time) for the impacts on Jupiter: A = 21, July 16.826 UT; B = 20, 17.113; C = 19, 17.287; D = 18, 17.483; E = 17, 17.625; F = 16, 18.014; G = 15, 18.308; H = 14, 18.805; K = 12, 19.425; L = 11, 19.919; N = 9, 20.428; P = 8 (= P_{2} = 8b), 20.624; Q = 7 (= Q_{1} = 7a), 20.831; R = 6, 21.223; S = 5, 21.627; T = 4, 21.758; U = 3, 21.907; V = 2, 22.166; W = 1, 22.330. Comparison with computations by P. W. Chodas and D. K. Yeomans, Jet Propulsion Laboratory, suggests that the uncertainty is now ∼ ±0.005 day for almost all these nuclei, the remaining uncertainty being mainly because most of the earlier observations were reduced using the STScI Guide Star Catalogue. Using the tidal-disruption model of Z. Sekanina, Chodas and Yeomans provide less certain impact times for lost and less well observed fragments: J = 13, July 19.11 UT; M = 10, 20.24; P_{1} = 8a, 20.69; Q_{2} = 7b, 20.81. They add that a significant number of particles beyond the eastern end of the nuclear train should now already have begun to strike Jupiter.

COMET McNAUGHT-RUSSELL (1993v) Ephemeris continuation from the orbital elements on MPC 23651:

1994 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 7 15^{h}11.^{m}37 +54^{o}28.^{′}7 1.579 1.852 88.^{o}3 33.^{o}3 13.2
17 15 27.83 +49 55.6 1.718 1.981 89.0 30.9 13.6
27 15 43.04 +45 32.2 1.863 2.110 89.2 28.8 14.1
Aug. 6 15 57.57 +41 21.3 2.017 2.237 88.7 27.0 14.5
16 16 11.75 +37 25.5 2.179 2.363 87.5 25.4 14.9
26 16 25.73 +33 47.1 2.350 2.487 85.7 23.9 15.3
Sept. 5 16 39.61 +30 27.2 2.529 2.610 83.2 22.6 15.7

Total visual magnitude estimates: Apr. 1.15 UT, 6.2 (C. E. Spratt, Victoria, BC, 11 × 80 binoculars); 15.93, 7.6 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector); May 2.07, 7.4 (J. E. Bortle, Stormville, NY, 10×50 binoculars); 15.88, 9.5 (L. Szentaskó, Veresegyház, Hungary, 0.33-m reflector); June 9.34, 11.1 (A. Hale, near Orogrande, NM, 0.41-m reflector).

1994 July 9 Brian G. Marsden